# Pentagonal nanowires from topological crystalline insulators: a platform for intrinsic core-shell nanowires and higher-order topology

### v1

@@ III.2 Electronic properties of the pentagonal nanowires
To see the effect of the thickness dependence on the electronic properties of pentagonal NWs, we considered the [110] oriented pentagonal SnTe NWs for different thicknesses. We report the electronic band structures including the SOC effects for two cases of the [110] oriented SnTe pentagonal NWs with Sn on the twin boundaries and Te on the twin boundaries. The band structures of the SnTe and PbTe pentagonal NWs without SOC included are instead shown in Appendix D, no major differences are observed between SnTe and PbTe. However, the two pentagonal NWs with different twin boundaries show different properties as observed also in [111] thin filmsSamadi et al. (2023). In Fig. 5 we show the two cases with different twin boundaries, where all the band structures reveal metallic behavior with bands connecting the valence and conduction electrons. While the behavior as a function of the thickness does not change for the Te twin-boundary NWs, for the 5-ring NW with Sn twin-boundary we observe an additional band connecting the valence and conduction band that seems to close and invert the band gap at the point. We have a reduction of the band gap at but not at the Z point as a strong indication that the two high-symmetry k-points are different from the topological point of view. In Appendix E, we show that the core chain (CC) of atoms is responsible for the metallicity in the systems. Indeed, the band connecting the valence and conduction electrons disappears when the core chain of atoms is removed within DFT calculations.
@@
In addition to the C symmetry of the global system, it was shown that also the local C symmetry is relevant. Since the preservation of local C symmetry in the region among the twinning is crucial for the observation of the higher-order topologyGeier et al. (2021), we study the band structure of the pentagonal NWs with [110] orientation and perfect local C symmetry (without structural relaxation). We define this crystal structure without structural relaxations as a high-symmetric pentagonal NW. We show the electronic band structures and charge densities for the most symmetric 5-ring SnTe pentagonal NW with Sn and Te on the twin boundaries in Fig. 6 and Fig. 7. In the band structure of Fig. 6(a) and Fig. 7(a) we observe a two-times degenerate band connecting the valence and conduction band with a positive slope for the Sn-twin boundary and negative slope for the Te-twin boundary. In the case of the Te-twin boundary, we calculate the charge density at the point associated with the band connecting valence and conduction in Fig. 7(b). Its charge is shared between the CC and the Sn atoms on the 5 mirror planes. Supported by the results presented in the model Hamiltonian, we assume that a similar state is present also in the Sn twin boundary NWs but in the conduction band and more hybridized with other bands.
@@
To gain further insight, we perform tight-binding calculations using the model Hamiltonian for a section of the SnTe pentagonal NWs. The tight binding model was reported in previous literatureBrzezicki et al. (2019); Nguyen et al. (2022) and it was derived from the Shiozaki-Sato-Gomi (SSG) modelShiozaki et al. (2015). We use as a section the mirror plane shown in Fig. 8(a) for a minimal model of the experimental system where the CC acts as a termination defect of the SSG model. In Fig. 8(b), the results for the Sn twin boundary describe a metallic band connecting the valence and conduction band associated with the CC. Even if this band is quite localized in the CC, the hybridization with the shell of the NWs described by the hopping t is extremely relevant. Indeed, without the coupling between core and shell, the metallic band disappears as we can see in Fig. 8(c). Exchanging anion and cation to change the twin boundary, we obtain a band structure for the Te twin-boundary shown in Fig. 8(d). With the Te twin boundary, we obtain the CC state to have a negative slope. Fig. 8(e) describes a simpler model with periodicity along the z-direction that further confirms the relevance of the CC as a termination defect of the SSG model for the formation of the metallic bands in pentagonal NWs. Basically, in both twin boundary cases, the CC band connects the valence and conduction bands enforcing the metallicity even in the ultrathin limit. Therefore, the pentagonal NWs are intrinsic core-shell NWs composed of a single material with a metallic chain in the core and an insulating cubic SnTe shell. Core and shell are electronically and structurally different but have the same constituent elements and stoichiometry.
III.3 Topological electronic states in the core chain of the pentagonal nanowires
@@
The electronic properties of the core chain were understood using a simplified SSG model Hamiltonian derived from a ribbon of a topologically non-trivial model with nonsymmorphic chiral Z invariantBrzezicki and Hyart (2020). This model (named also the Rice-Mele model) was used also to study the presence of topological midgap corner modes in n-fold rotational symmetric chiral insulators to investigate the existence of a HOTI protected by rotational symmetryvan Miert and Ortix (2020). The cross-section of the NW was used for the model Hamiltonian is shown in Fig. 8(a) with the three hopping parameters of the model names as t, t and t. When the hopping t is zero, there are no bands connecting the valence and conduction band and the system is an insulator as shown in Fig. 8(c). When t is different from zero, this model gives electronic states inside the gap connecting the valence and conduction bands resulting in a symmetry-enforced metallic phase. In the case of Sn on the twin boundary shown in Fig. 8(b), the connecting band has a positive slope while has a negative slope in the case of Te on the twin boundary plotted in Fig. 8(d). The robustness of this symmetry-enforced metallic phase was confirmed also by linear scaling DFT calculations (Appendix F) and by DFT calculations of the NWs under isotropic pressure (Appendix G). We understand how the electronic properties critically depend on the composition of the CC. The minimal model to reproduce this metallic phase is an SSG model with a termination defect as shown in Fig. 8(e). Note that the pure SSG model without termination defect exhibits no end-states Brzezicki and Hyart (2020). This is because the termination of the system breaks the nonsymmorphic chiral symmetry that protects the Z topological invariant. However, in the presence of the topological invariant, there is charge accumulation at the edges, in strict analogy to the 1D systems with inversion symmetryRhim et al. (2017). We find that with termination defect such accumulation can again lead to the appearance of the end-states.
@@
III.4 Signature of fractional charge and higher-order topology in pentagonal nanowires
@@
Now, we move to describe the signatures of higher-order topology as the hinge states. The hinge states for SnTe cubic nanowires (grown along the [001] orientation presenting a square section) were observed at a given k-point that was close to the edge of the Brillouin zoneNguyen et al. (2022); Skiff et al. (2023). In the case of the Sn twin boundary, we analyzed the charge density at the point in 6(b). Localized charge densities are visible at the corners of the 5-ring pentagonal NW. We surmise that this state will evolve in the localized higher-order topological state once the thickness of the shell is large enough to generate the TCI state of SnTe. While recent studies focused on the higher-order topology in finite systems in the xy plane, more studies are necessary for the higher-order topology in NWs that are periodic along the z-direction and finite in the xy plane.
@@
In addition to hinge states, another feature of the higher-order topology is the presence of a fractional charge at the center. Fig. 9(a) and 9(b) show the charge for the SnTe cubic and pentagonal NWs, respectively. We report for comparison also the charge in the bulk SnTe. The charges on both Sn and Te of the CC substantially increase with respect to the charge in the bulk and are constant as a function of the thickness. This excess of charge could be associated with the fractional charge appearing in the presence of disclinations. The fractional charge can be observed with and without topology. The fractional charge and higher-order topology are strongly developed in 2D materials but it less developed for NWs. We leave it to future studies for a deeper explanation of the fractional charge. The interplay between the topological properties from the core and from the shell will be described elsewhere.Samadi and Buczko (2024)
@@ IV Conclusions
We have synthesized ionic pentagonal NWs of PbSnTe with smooth texture devoid of imperfections. Within a semiclassical approach, the pentagonal crystal structure is stabilized by an interplay between the ionic bond and the covalent bond. Within a full quantum approach, we have demonstrated that the pentagonal phase of the SnTe NWs is a metastable phase close in energy to the ground state which is the cubic structure. The pentagonal NWs are an intrinsic core-shell system where the core and shell have different structural, electronic and topological properties. The pentagonal NWs with [110] orientations are metallic due to the presence of a single band connecting the conduction and valence bands coming from the one-dimensional disclination in the core of the NW. This connecting band clearly comes from the one-dimensional disclination and it is related to the properties of the topological SSG model. We have found an excess charge in the center that could be associated with a fractional charge. We have calculated the charge in the real space related to the degenerate electronic states at the point and we find that these states are associated with higher-order topological states at the hinges of the NWs. The pentagonal NWs are a platform for higher-order topology and fractional charge. These pentagonal NWs represent a unique case of intrinsic core-shell NWs with distinct structural, electronic and topological properties between the core and the shell.
Acknowledgements.
We thank T. Dietl, C. Ortix, R. Buczko and S. Samadi for useful discussions. The work is supported by the Foundation for Polish Science through the International Research Agendas program co-financed by the European Union within the Smart Growth Operational Programme (Grant No. MAB/2017/1). W.B. acknowledges support from Narodowe Centrum Nauki (NCN, National Science Centre, Poland) Project No. 2019/34/E/ST3/00404. J.S. and S.K. acknowledge support from the National Science Centre Poland through the projects 2019/35/B/ST3/03381 and 2019/35/B/ST5/03434. C.M.C. and S.S. thank Carl Tryggers Stiftelsen (CTS 20:71) and the Swedish Research Council (VR 2021-04622) for financial support. We acknowledge the access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant g91-1418, g91-1419 and g91-1426 for the availability of high-performance computing resources and support. We acknowledge the CINECA award under the ISCRA initiative IsC99 ”SILENTS”, IsC105 ”SILENTSG”, IsB26 ”SHINY” and IsB27 ”SLAM” grants for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center Grant No. 609 (pl0223-01) and pl0267-01. The computations at Linnaeus University were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Dardel partially funded by the Swedish Research Council through grant agreement no. 2022-06725.Appendix A Experimental setup
PbSnTe NWs were grown by molecular beam epitaxy technique in a home-built system equipped with SnTe, Pb and Te solid sources. We used commercial (Azelis, France) both (100)- and (111)-oriented p-Si substrates covered with a relatively thin oxide layer (up to 10 Å) such that we observed slight streaky reflection high-energy electron diffraction (RHEED) patterns even without any surface preparation. Covering the surface with an amorphous SiO layer prevents deposition directly on monocrystalline silicon. On the other hand, these oxides support the mobility of the adatoms on the surface. To uncover Si, the etching procedure in HF was employed. Different solutions of acid (from 8% to 40%), as well as various times of etching (up to 2 min), were applied resulting in different oxides covering stages. The aim was to achieve separated pinholes in oxide which can play the role of the base for the nucleation center. Before loading into load-lock the freshly etched Si substrates were kept in ethanol to protect them against the air. Instead of the typically exploited vapor-liquid-solid (VLS) mode of growth effective for IV-VI, e.g. using Au nanodroplets as a catalyst Dziawa et al. (2010), we applied vapor-solid (VS) mode. The growth temperature was in the range from 450 °C to 500 °C. The SnTe/Pb molecular beam flux ratio was kept to obtain in PbSnTe. The total molecular flux used to grow NWs corresponds to the rate of growth for layers from 0.1 Å/s to 0.2 Å/s. This results in NWs longer than 1 m and the rate of NWs growth higher than 0.5 m/h. Diameters of the NWs show relatively wide distribution from the smallest around 15 nm to even 200 nm with the maximum at approximately 40-80 nm. Under these conditions both types of NWs are observed, i.e., exhibiting four-fold and five-fold symmetry of cross-section (see Fig. 1a).
SEM observations were made using Scanning Electron Microscope ZEISS Auriga - CrossBeam Workstation, images were made at 5 kV with magnification from 10k to 300k x. The images show that the length of the NWs varies from 0.7 m to 3 m, and the width of the structures ranges between 40-150 nm. The surface density is from 0.1 to even 5 m. On the substrate, we observed five-fold NWs with greater amounts, less often four-fold NWs. Structures are symmetrical and well-shaped with visible blocks. The surface of the NWs was smooth without any imperfections. Structural properties of the PbSnTe NWs were examined using FEI-Titan 80-300 transmission electron microscope operating at 300 kV, equipped with an image corrector and HAADF scanning transmission electron microscopy detector. The cross-sectioned NWs were prepared with the Thermo Scientific’s Helios G1 NanoLab DualBeam 600 Scanning Electron Microscope, with FIB (Focused Ion Beam)- resolution of 5.0 nm at 30 kV, and lift-off procedure for cross-sections with an Omni-Probe nanomanipulator. The cross-sectioned NWs were investigated in the zone axis [110] which is parallel to the NWs growth axis, in scanning transmission electron microscopy as well as in high-resolution transmission electron microscopy mode. The five-fold symmetry was confirmed, and the (111) planes twin boundaries were visible as well as brighter areas on the twin. These areas are related to strain fields that are visible to distinguish for observations under camera length 0.230 m. EDS determined the averaged elemental ratio as Sn:Pb = 55:45 (1:1).
Appendix B Computational details
Electronic structure calculations were performed within the framework of the first-principles density functional theory based on plane wave basis set and projector augmented wave method using VASPKresse and Furthmüller (1996) package. The structural relaxations are performed without SOC, while all the band structure calculations of the main text are fully relativistic by considering SOC, a plane-wave energy cut-off of 250 eV has been used. As an exchange-correlation functional, the generalised gradient approximation of Perdrew, Burke, and Ernzerhof has been adoptedPerdew et al. (1996). We have performed the calculations using 1112 k-points centered in . The
electronic band structures are computed with 60 k-points between the and Z direction.
The convergence criterion for the force is considered to be 0.01 eV/Å, while
eV of energy tolerance is taken for structural relaxation.
In the high-symmetric structure, we study the systems without structural relaxation to study the system with perfect C symmetry in the regions among the twinnings.
Appendix C Semiclassical approach to describe the stability of ionic pentagonal NWs
Since the pentagonal geometry is poorly studied, we also calculated some semiclassical quantities to unveil the reasons for its structural stability. For instance, the Madelung constant and the particle density for three types of polygons such as cube, pentagon, and hexagon are calculated. Finally, we combine both the properties in one plot showing that the pentagonal NWs can be formed in case of an interplay between ionic and covalent bonds.
The three types of polygons considered are shown in Fig. 10. Two types of atoms with different atomic radii are studied and general equations for particle density in two dimensions (2D) are used. The atoms are positioned on the sides and corners of the polygon, the size of the side atoms (r) is variable and smaller than the corner atoms (R). Hence, two regions are studied i.e. and . Here is the critical value of the radius for which the side atom touches the corner one and the density becomes maximum. The general formula for the particle density D in the case of NWs with a polygonal section having sides (N is 4, 5 and 6 for the square, pentagon and hexagon respectively) can be written as;
We define x=r/R where r is the radius of the side atoms and R is the radius for the atom at the corners. The particle density D(x) reaches a maximum at a critical value x=x and afterward starts to decrease. In the cubic environment, the particle density is calculated to be 0.785 for = 0 ( = 0), while it is increased to 0.919 if the radius of the atom located on the side of the polygon approaches the critical value by touching the corner atoms. However, the density again decreases for and becomes 0.785 when . An almost similar trend is followed by pentagons and hexagons; Following these equations, the particle densities at = 0 are reported in Table 1. We got the highest density for the hexagonal structure, lowest for the cubic, and intermediate for the pentagon when . The hexagonal structure possesses the highest particle density revealing that the covalent bond is favorable in this case, which means more covalent character pushes the system to stay in the hexagonal phase.
Polygon | Particle density | Madelung constant |
Cube | 0.785 | 1.172 |
Pentagon | 0.825 | 0.694 |
Hexagon | 0.906 | 0.464 |
The Madelung constant determines the electrostatic potential for a single ion in a crystal by approximating the ions with point charges. We calculate the Madelung constant for cube, pentagon, and hexagon. Generally, the Madelung constant is defined as:
where is the number of j-th ions, is the distance between the ions with N= 4, 5, and 6 for the cubic, pentagonal, and hexagonal structures, respectively. Considering (opposite for the two atoms) as the charge and as the distance between the two atoms, in the case of a cubic structure, we have:
In the case of a pentagonal structure:
while for the hexagonal structure:
The value of the Madelung constant is very crucial in the case of ionic bonds. As expected, we got the highest value of the Madelung constant for the cubic system (1.172) and the lowest (0.464) for the hexagon, while the pentagon bears the intermediate value (0.694). The large value of the Madelung constant for a cubic system indicates the highest contribution to ionicity and implies an ionic insulator scenario, which favors the system to be in the cubic phase.
This manifestation of the highest Madelung constant and lowest particle density for the cube indicates the dominance of ionic character in the cubic phase, while the lowest Madelung constant and highest particle density for the hexagonal structure indicate that the system presents a more covalent character in this phase.
On the other hand, intermediate values of Madelung constant and particle density for the pentagonal shape (see Fig. 11) reveal the contribution of both the ionic and covalent characters in the bond formation, which implies that if the bond is partially ionic and partially covalent, the pentagonal phase is expected to be the most stable. In the NWs based on lighter elements such as Se and Sn, the bonds would be more ionic, therefore we expect to be less likely to form a pentagonal crystal structure. Therefore, NWs based on Pb and Te would be more likely to form pentagonal NWs.
To summarize, from a semiclassical point of view, we understand that the interplay between covalent and ionic bonds favors the formation of pentagonal NWs. The quantum calculations reported in Section III of the main text demonstrated that the pentagonal NWs are metastable with the ground state being the cubic NWs. The difference between the ground state energies of cubic and pentagonal NWs becomes very small for thick NWs paving the way for the experimental formation of pentagonal NWs.
Appendix D Band structure of the [110] pentagonal SnTe and PbTe NWs without SOC
In Fig. 12 we show the band structures of these systems without considering SOC interaction; here, both the compounds manifest bands that connect valence and conduction regions except 1-ring PbTe NW, which has an indirect band gap of 0.423 eV. The band connecting valence and conduction at the point is in the conduction band and goes in the valence band increasing as opposite to the pentagonal NW with [001] orientation. In Appendix E, we show that the core chain is responsible for the metallicity in the systems. Indeed, the band connecting the valence and conduction electrons disappears when the core chain of atoms is removed. Therefore, this band connecting valence and conduction is not related to topological properties since is not a surface band. This band, coming from the one-dimensional core chain, is composed of p orbitals of Sn and Te with a very short distance, therefore it would be very difficult to reproduce within the tight-binding models based on a bulk cubic environment with the normal lattice constant. The central atoms (a Sn and Te atom in the unit cell) have around 0.5 electrons more than other atoms, while the first and second-ring atoms have a little less charge of around 0.01 electrons. These 0.5 electrons are due to the different structural environments of the CC or to the fractional charge.
Appendix E Removing the core chain
Here, we calculate the band decomposed charge density at different k-points along the direction for the SnTe pentagonal NWs oriented along [110]. To check the influence of the core chain on the electronic structure, we show the band decomposed charge density in Fig. 13. The charge density for the band connecting the valence and conduction sectors is distributed in the center of NW in real space when we consider a point in the middle of this band, namely not hybridized. Instead, when we take points where this band is hybridized with other electronic bands, the charge density is distributed in other regions as well, as shown in Fig. 13a). These results clearly indicate that the core chain is responsible for the metallicity in the systems. Further confirming this claim, we removed the core chain of the 3-ring pentagon (the remaining crystal structure is composed of only atoms in a NaCl-phase environment with 6 nearest neighbors) and calculated its band structure as illustrated in Fig. 13b). The band connecting the valence and conduction electrons disappears when the core chain of atoms is removed. The pentagonal SnTe NW without the core chain becomes a simple band insulator as expected for SnTe NWs with a cubic environment in the ultrathin limit. We observe the same behavior also for the [001] pentagonal NWs.
Appendix F Reproducing the electronic structure within the ONETEP code
To double-check the results obtained within VASP, we perform electronic structure calculations with a DFT code that uses a different basis for the charge density as ONETEPSkylaris et al. (2005); Prentice et al. (2020). The computational framework was presented in a previous paperHussain et al. (2023). The results are reported in Figs. 14 and 15 for the 5-rings optimized structure of SnTe with Sn and Te twin boundaries, respectively. The results for the 5-rings of SnTe with high-symmetry structure and with Sn and Te twin boundaries are reported in Figs. 16 and 17, respectively. In order to do this, we repeat the unit cell three times along the c-axis, this produces a triple downfolding of the band structure. Due to the downfolding, the metallic band has an opposite inclination with respect to the VASP but the results are strongly consistent with the VASP results. The SnTe band structures within ONETEP show a slightly larger gap for the shell bands as it happens in the cubic NWsHussain et al. (2023).
Appendix G Band structure under applied pressure
The study of the topological properties in the cubic NWs is extremely demanding computationally because of the thickness dependence of the topology. Indeed, we need to have a large thickness to reach the topological phase in the cubic NWs. In cubic bulk SnTe and PbTe compounds, the effect of the compression pushes the SnTe material-class compounds toward the topological region of the phase diagram. Since our thin NWs are extremely thin the cubic shell region is in a trivial regime, we apply compression in order to observe the topological transition in the shell region. The transition of the cubic shell from the trivial to the topological phase for the pentagonal NWs should go through a closure of the trivial band gap at one or both high-symmetry points and Z. The effect of compression on the electronic band structures of symmetric 5-ring SnTe pentagonal NW with Te on the twin boundaries is shown in Fig. 18. We notice a reduction of the band gaps at and Z, the systems become more metallic as the compression is increased due to the increasing of the bandwidth as happens in bulk phasesHsieh et al. (2012), however, no closure of the gap at or Z is observed. Even if the thickness is too thin to develop a topological phase in the cubic SnTe, we have the connection between valence and conduction due to the core chain persisting at every pressure.

### v3

@@ III.2 Electronic properties of the pentagonal nanowires
To see the effect of the thickness dependence on the electronic properties of pentagonal NWs, we considered the [110] oriented pentagonal SnTe NWs for different thicknesses. We report the electronic band structures including the SOC effects for two cases of the [110] oriented SnTe pentagonal NWs with Sn on the twin boundaries and Te on the twin boundaries. The band structures of the SnTe and PbTe pentagonal NWs without SOC included are instead shown in Appendix D in the Supp. Information, no major differences are observed between SnTe and PbTe. However, the two pentagonal NWs with different twin boundaries show different properties as observed also in [111] thin filmsSamadi et al. (2023). In Fig. 5 we show the two cases with different twin boundaries, where all the band structures reveal metallic behavior with bands connecting the valence and conduction electrons. While the behavior as a function of the thickness does not change for the Te twin-boundary NWs, for the 5-ring NW with Sn twin-boundary we observe an additional band connecting the valence and conduction band that seems to close and invert the band gap at the point. We have a reduction of the band gap at but not at the Z point as a strong indication that the two high-symmetry k-points are different from the topological point of view. In Appendix E in the Supp. Information, we show that the core chain (CC) of atoms is responsible for the metallicity in the systems. Indeed, the band connecting the valence and conduction electrons disappears when the core chain of atoms is removed within DFT calculations.
@@
In addition to the C5 symmetry of the global system, it was shown that also the local C4 symmetry is relevant. Since the preservation of local C4 symmetry in the region among the twinning is crucial for the observation of the higher-order topologyGeier et al. (2021), we study the band structure of the pentagonal NWs with [110] orientation and perfect local C4 symmetry (without structural relaxation). We define this crystal structure without structural relaxations as a high-symmetric pentagonal NW. We show the electronic band structures and charge densities for the most symmetric 5-ring SnTe pentagonal NW with Sn and Te on the twin boundaries in Fig. 6 and Fig. 7. In the band structure of Fig. 6(a) and Fig. 7(a) we observe a two-times degenerate band connecting the valence and conduction band with a positive slope for the Sn-twin boundary and negative slope for the Te-twin boundary. In the case of the Te-twin boundary, we calculate the charge density at the point associated with the band connecting valence and conduction in Fig. 7(b). Its charge is shared between the CC and the Sn atoms on the 5 mirror planes. Supported by the results presented in the model Hamiltonian, we assume that a similar state is present also in the Sn twin boundary NWs but in the conduction band and more hybridized with other bands.
@@
To gain further insight, we perform tight-binding calculations using the model Hamiltonian for a section of the SnTe pentagonal NWs. The tight binding model was reported in previous literatureBrzezicki et al. (2019); Nguyen et al. (2022) and it was derived from the Shiozaki-Sato-Gomi (SSG) modelShiozaki et al. (2015) (for more see Appendix H in Supp. Information). We use as a section the mirror plane shown in Fig. 8(a) for a minimal model of the experimental system where the CC acts as a termination defect of the SSG model. In Fig. 8(b), the results for the Sn twin boundary describe a metallic band connecting the valence and conduction band associated with the CC. Even if this band is quite localized in the CC, the hybridization with the shell of the NWs described by the hopping t3 is extremely relevant. Indeed, without the coupling between core and shell, the metallic band disappears as we can see in Fig. 8(c). Exchanging anion and cation to change the twin boundary, we obtain a band structure for the Te twin-boundary shown in Fig. 8(d). With the Te twin boundary, we obtain the CC state to have a negative slope. Fig. 8(e) describes a simpler model with periodicity along the z-direction that further confirms the relevance of the CC as a termination defect of the SSG model for the formation of the metallic bands in pentagonal NWs. Basically, in both twin boundary cases, the CC band connects the valence and conduction bands enforcing the metallicity even in the ultrathin limit. Therefore, the pentagonal NWs are intrinsic core-shell NWs composed of a single material with a metallic chain in the core and an insulating cubic SnTe shell. Core and shell are electronically and structurally different but have the same constituent elements and stoichiometry.
III.3 Tunability of pentagonal nanowires growth
Starting from the general aspects, we need to weaken the ionic bonds to favor the formation of pentagonal nanowires as we discussed in the previous subsections. Once the ionic part of the bond is weakened and the system is more covalent, the formation of pentagonal structures depends on the ratio of growth rates in the (100) and (111) directions, which is responsible for the twinningWild et al. (1994); Bühler and Prior (2000). Many works on fivefold symmetry focus on nanoparticles. In addition to elemental materials, various covalent alloys, compounds, and composite nanoparticles grow with fivefold twinned structuresHofmeister (2004)
Regarding the specific (Pb,Sn)Te system, we report what are the growth conditions that favor the formation of the pentagonal nanowires. We have hints of nucleation, but no conclusive evidence of nucleation of the central chain from several experiments (TEM of the NW//Si interface; STM at the center of the cross-section; detailed analysis of growth as a function of conditions). In the case of ternary (Pb,Sn)Te, the deviation of the chemical composition from x=0.5 (estimated for the layers based on the base of beam equivalent pressure) dramatically reduces the occurrence of 5-fold NWs. A change in the Te/Pb flux ratio of about 20% from the optimal for 5-fold NWs growth results in growth NWs with 4-fold symmetry. Therefore, in order to observe the pentagonal nanowires the doping concentration should be around x=0.50 i.e. close to the critical concentration where there is the topological transition at room temperatureDornhaus et al. (1983). Based on the experimental data, the slight excess of Pb in the total flux of (Pb,Sn)Te supports the nucleation of islands with other than a 4-fold structure, this agrees with our theoretical results supporting the presence of heavy atoms as necessary for the pentagonal NWs. Regarding the temperature, for similar values of SnTe, Pb, and Te molecular fluxes, NWs 5-fold are observed at higher growth temperatures in the range where cubic NWs growth is not observed at all. Therefore, a high temperature is required since our theoretical results suggest that the pentagonal structure is metastable at low temperatures.
Additionally, the crystal structure can be switched from cubic to pentagonal during the growth process to create heterostructure nanowires as shown in Fig. 9. In particular, this is the result of the PbTe/(Pb,Sn)Te//Si(100) growth process. In the first part of the growth, the growth time of (Pb,Sn)Te was performed, while in the second part, we grew PbTe. The growth of PbTe was carried out by turning off the SnTe flux for initial (Pb,Sn)Te and the PbTe flux remained unchanged. Hence, it suggests that both parts consist of (Pb,Sn)Te rather than PbTe/(Pb,Sn)Te. Further studies would be needed to establish the chemical composition of both parts of these NW heterostructures. More technical details are reported in Appendix A in the Supp. Information.
III.4 Topological electronic states in the core chain of the pentagonal nanowires
@@
The electronic properties of the core chain were understood using a simplified SSG model Hamiltonian derived from a ribbon of a topologically non-trivial model with nonsymmorphic chiral Z2 invariantBrzezicki and Hyart (2020). This model (known also as the Rice-Mele model) was used also to study the presence of topological midgap corner modes in n-fold rotational symmetric chiral insulators to investigate the existence of a HOTI protected by rotational symmetryvan Miert and Ortix (2020). The cross-section of the NW that was used for the model Hamiltonian is shown in Fig. 8(a), with the three hopping parameters of the model named as t1, t2 and t3. When the hopping t3 is zero, there are no bands connecting the valence and conduction band and the system is an insulator as shown in Fig. 8(c). When t3 is different from zero, this model gives electronic states inside the gap connecting the valence and conduction bands resulting in a symmetry-enforced metallic phase. In the case of Sn on the twin boundary shown in Fig. 8(b), the connecting band has a positive slope while has a negative slope in the case of Te on the twin boundary plotted in Fig. 8(d). The robustness of this symmetry-enforced metallic phase was confirmed also by linear scaling DFT calculations (Appendix F in the Supp. Information) and by DFT calculations of the NWs under isotropic pressure (Appendix G in the Supp. Information). These results demonstrate that the electronic properties critically depend on the composition of the CC. The minimal model that reproduces this metallic phase is a SSG model with a termination defect as shown in Fig. 8(e). Note that a pure SSG model without termination defects exhibits no end-states Brzezicki and Hyart (2020). This is because the termination of the system breaks the nonsymmorphic chiral symmetry that protects the Z2 topological invariant. However, in the presence of the topological invariant, there is charge accumulation at the edges, in strict analogy to the 1D systems with inversion symmetryRhim et al. (2017). We find that with termination defect such accumulation can again lead to the appearance of the end-states.
@@
III.5 Signature of fractional charge and higher-order topology in pentagonal nanowires
@@
Now, we move to describe the signatures of higher-order topology as the hinge states. The hinge states for SnTe cubic nanowires (grown along the [001] orientation presenting a square section) were observed at a given k-point that was close to the edge of the Brillouin zoneNguyen et al. (2022); Skiff et al. (2023). In the case of the Sn twin boundary, we analyzed the charge density at the point in 6(b). Localized charge densities are visible at the corners of the 5-ring pentagonal NW. We surmise that this state will evolve in the localized higher-order topological state once the thickness of the shell is large enough to generate the TCI state of SnTe. While there exist recent studies focused on higher-order topology in finite systems in the xy plane, additional studies are necessary for the higher-order topology in NWs that are periodic along the z-direction and finite in the xy plane.
@@
In addition to hinge states, another feature of the higher-order topology is the presence of a fractional charge at the center. Fig. 10(a) and 10(b) show the charge for the SnTe cubic and pentagonal NWs, respectively. We report for comparison also the charge in the bulk SnTe. The charges on both Sn and Te of the CC substantially increase with respect to the charge in the bulk and are constant as a function of the thickness. This excess of charge could be associated with the fractional charge appearing in the presence of disclinations. The fractional charge can be observed with and without topology. Fractional charge and higher-order topology are strongly developed in 2D materials but there are less evident in NWs. We leave it to future studies to provide for a deeper explanation of the fractional charge. The interplay between the topological properties from the core and from the shell will be described elsewhere.Samadi and Buczko (2024)
@@ IV Conclusions
We have synthesized ionic pentagonal NWs of Pb1-xSnxTe with smooth texture devoid of imperfections. Within a semiclassical approach, the pentagonal crystal structure is stabilized by an interplay between the ionic bond and the covalent bond. Within a full quantum approach, we have demonstrated that the pentagonal structure of the SnTe NWs is a metastable phase at T=0 but close in energy to the ground state which is the cubic structure. Therefore, we need to weaken the ionic bond with the presence of heavy atoms Pb and Te and to be close to the topological transition, finally, the growth of the pentagonal NWs should take place at high temperatures since they are not stable at low temperatures. The pentagonal NWs are an intrinsic core-shell system where the core and shell have different structural, electronic and topological properties. The pentagonal NWs with [110] orientations are metallic due to the presence of a single band connecting the conduction and valence bands coming from the one-dimensional disclination in the core of the NW. This connecting band clearly comes from the one-dimensional disclination and it is related to the properties of the topological SSG model. We have found an excess charge in the center that could be associated with a fractional charge. We have calculated the charge in the real space related to the degenerate electronic states at the point and we find that these states are associated with higher-order topological states at the hinges of the NWs. The pentagonal NWs are a platform for higher-order topology and fractional charge. These pentagonal NWs represent a unique case of intrinsic core-shell NWs with distinct structural, electronic and topological properties between the core and the shell.
Author contributions
G. Hussain: Investigation, Methodology, Validation, Data curation, Writing - original draft. G. Cuono: Investigation, Methodology, Validation, Data curation, Writing - original draft. P. Dziawa: Investigation, Validation, Data curation, Writing - original draft. D. Janaszko: Investigation, Validation, Data curation, Writing - original draft. J. Sadowski: Funding acquisition. S. Kret: Funding acquisition. B. Kurowska: Investigation, Methodology, Validation, J. Polaczyński: Conceptualization, Investigation, Methodology, Validation, K. Warda: Investigation, Methodology, Validation, Data curation, Writing - original draft. S. Sattar: Writing - review & editing. C. M. Canali: Funding acquisition, Resources, Writing - review & editing. A. Lau: Validation W. Brzezicki: Methodology, Investigation, Validation, Data curation, Funding acquisition. T. Story: Supervision, Resources. Carmine Autieri: Supervision, Conceptualization, Project administration, Resources, Writing - review & editing.
Conflicts of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
We thank T. Dietl, C. Ortix, R. Buczko and S. Samadi for useful discussions.
The work is supported by the Foundation for Polish Science through the International Research Agendas program co-financed by the European Union within the Smart Growth Operational Programme (Grant No. MAB/2017/1).
W.B. acknowledges support from Narodowe Centrum
Nauki (NCN, National Science Centre, Poland) Project No. 2019/34/E/ST3/00404.
J.S. and S.K. acknowledge support from the National Science Centre Poland through the projects 2019/35/B/ST3/03381 and 2019/35/B/ST5/03434. C.M.C. and S.S. thank Carl Tryggers Stiftelsen (CTS 20:71) and the Swedish Research Council (VR 2021-04622) for financial support.
We acknowledge the access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant g91-1418, g91-1419 and g91-1426 for the availability of high-performance computing resources and support.
We acknowledge the CINECA award under the ISCRA initiative IsC99 ”SILENTS”, IsC105 ”SILENTSG”, IsB26 ”SHINY” and IsB27 ”SLAM” grants for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center Grant No. 609 (pl0223-01) and pl0267-01. The computations at Linnaeus University were enabled by resources provided by the National Academic Infrastructure for Supercomputing
in Sweden (NAISS) at Dardel partially funded by the Swedish Research Council through grant agreement no. 2022-06725.